# Regularized Exponentially Tilted Empirical Likelihood for Bayesian Inference

### 28 Dec 2023 v1

@@ 1 Introduction
However, both EL and ETEL have an inherent limitation in that they are only defined on a proper subset of the original parameter space due to the convex hull constraint or empty set problem (Grendár & Judge, 2009). By convention, the likelihoods are set to zero for parameter values that violate the convex hull constraint. For Bayesian inference, the zeroes in the likelihood imply a restricted posterior domain. This is conceptually unsatisfactory as, with a larger sample size, the convex hull may expand and the likelihood become positive. Additionally, as the restricted domain is often non-convex, (Chaudhuri et al., 2017), a more sophisticated posterior sampling scheme may be needed to fit the model.
@@
In this paper, we propose a method to address the convex hull constraint for Bayesian ETEL. While previous proposals have primarily focused on EL and frequentist inference, our proposal builds upon the AEL framework, introducing notable distinctions. First, we extend the method to accommodate fractional observations, following the approach of (Hainmueller, 2012). Second, we allow for the incorporation of multiple pseudo observations. Third, we pass to the limit, ensuring that the convex hull constraint is satisfied for all parameter values simultaneously. This resulting formulation naturally induces a form of regularization that removes the constraint. Our method’s main contributions encompass: (i) addressing the convex hull constraint for ETEL while retaining desirable asymptotic properties; (ii) enhancing stability and robustness of small-sample performance compared to existing methods; (iii) providing flexibility in Bayesian modelling and allowing one to incorporate a novel form of prior information.
This paper is organized as follows. In Section 2, we introduce the notation used in the paper and provide a brief overview of ETEL. Then, we propose a weighted version of ETEL that incorporates fractional pseudo-data with the maximum entropy reweighting scheme. In Section 3, we propose inducing regularization on the formulation of ETEL, exploring two equivalent approaches: (i) a limiting procedure with fractional pseudo-data and (ii) direct incorporation of a continuous exponential family distribution in the minimization of the KL divergence. We derive asymptotic properties of the proposed methods. In Section 4, we evaluate the performance of the methods through simulation studies. In Section 5, we present an application to the estimation of median income for four-person families. Finally, we conclude with a discussion of directions for future research in Section 6. The proofs of the theoretical results are provided in the supplementary materials. The implementation is available as the R package retel, along with the simulation and data analysis code, in a GitHub repository at https://github.com/markean/retel.
@@ 4.1 Posterior Coverage
To investigate the validity of RETEL for Bayesian inference, we begin by simulating a value of ${\theta}$ from a logistic distribution denoted as ${\textnormal{Logistic}(l,s)}$, where ${l}$ is the location parameter and ${s}$ is the scale parameter. Next, we generate ${n}$ observations from ${N(\theta,1)}$ and compute ${H}$ for the two versions of RETEL. Throughout the analysis, we employ ${p_{n}(\theta,\lambda)}$ in Equation 4 with ${\mu_{n,\theta}}={\overline{X}-\theta}$ and ${\Sigma_{n,\theta}}={1}$ for the univariate mean parameter ${\theta}$. For comparison purposes, we also compute ${H}$ using ETEL and AETEL. Keeping ${l}$ fixed at 00{0}, we repeat this procedure $10000$ times for each combination of ${n}\in{\{5,20,50,100\}}$, ${s}\in{\{1,5\}}$, and ${\tau_{n}}\in{\{1,\log n\}}$. We approximate the posterior distributions on a grid of ${\theta}$ values. Using the computed ${H}$ values, we conduct the Kolmogorov–Smirnov test to evaluate the uniformity of the distributions. The resulting ${p}$-values are reported in Table 1.
@@
With a smaller sample size of ${n}={20}$, ${\textnormal{RETEL}_{f}}$ tends to show a closer conformity to ${U(0,1)}$ compared to ETEL and AETEL. The impact of a larger prior variance (${s}={5}$) and a larger ${p_{c}(\theta)}$ (${\tau_{n}}={\log n}$) becomes more apparent when ${n}={20}$. As the sample size increases, the differences between the posterior distributions of the methods become negligible. When ${n}={100}$, none of the methods exhibit a significant departure of ${H}$ from ${U(0,1)}$. In all settings, ${\textnormal{RETEL}_{f}}$ produces larger ${p}$-values than ${\textnormal{RETEL}_{r}}$. Figure 3 displays the quantile-quantile plots for the distribution of ${H}$ versus ${U(0,1)}$ when ${n}={5}$, ${s}={5}$, and ${\tau_{n}}={1}$. The plots highlight the differences in the tails of the distributions that are not apparent from the ${p}$-values alone. We emphasize that the Kolmogorov–Smirnov tests are based on a sample of $10000$ replicates and so are able to pick up quite small departures from the uniform distribution. All of the methods provide an excellent approximation to the null distribution when ${n}$ is 5050{50}50 or more. Additional plots for the full results are provided in Section 8 in the supplementary materials.
Figure 3: Quantile-quantile plots for the distribution of ${H}$ versus ${U(0,1)}$ when ${n}={5}$, ${s}={5}$, and ${\tau_{n}}={1}$. The light-tailed distribution from ETEL is due to the convex hull constraint. AETEL produces a heavier-tailed distribution than ${\textnormal{RETEL}_{f}}$.
@@
The results for ${l}={0}$ are presented in Table 2, where the prior mean matches the true parameter value. It can be seen that for all methods, as the sample size increases, the intervals become shorter and the coverage rates approach the target of ${95\%}$. As ${s}$ decreases, indicating stronger prior information on ${\theta}$ at 00{0}, higher coverage rates and shorter intervals are obtained. The differences between the methods are most pronounced when ${n}={5}$. The intervals obtained from ETEL exhibit significantly lower coverage rates compared to the other methods. AETEL produces the widest intervals with coverage rates higher than the nominal level. The wider intervals and departure from the nominal coverage rate are related to the boundedness problem of AEL, which arises due to the addition of one pseudo-observation (Emerson & Owen, 2009). On the other hand, RETEL yields higher coverage rates than the nominal level, but with much shorter intervals compared to AETEL. Within RETEL, ${\textnormal{RETEL}_{f}}$ produces wider intervals with higher coverage rates than ${\textnormal{RETEL}_{r}}$, consistent with the findings from the plots in Figure 3.
@@ 4.2 Expected Kullback–Leibler Divergence
    Repeat Steps 1–3 $10000$ times and take the average of the estimates from Step 3.
Step 2 is implemented with two chains of length $5000$, ensuring that the potential scale reduction factor (Gelman & Rubin, 1992) remains below 1.11.1{1.1}1.1 on average for each method. For the regularized methods, ${\tau_{n}}={1}$ is used when ${n=2}$, and ${\tau_{n}}={\log n}$ is used otherwise. We implement EL using the R package melt (Kim, 2023).

### 20 Feb 2024 v3

@@ 1 Introduction
However, both EL and ETEL have an inherent limitation in that they are only defined on a proper subset of the original parameter space due to the convex hull constraint or empty set problem (Grendár & Judge, 2009). By convention, the likelihoods are set to zero for parameter values that violate the convex hull constraint. For Bayesian inference, the zeroes in the likelihood imply a restricted posterior domain. This is conceptually unsatisfactory as, with a larger sample size, the convex hull may expand and the likelihood become positive. Additionally, as the restricted domain is often non-convex (Chaudhuri et al., 2017), a more sophisticated posterior sampling scheme may be needed to fit the model.
@@
In this paper, we propose a method to address the convex hull constraint for Bayesian ETEL. While previous proposals have primarily focused on EL and frequentist inference, our proposal builds upon the AEL framework, introducing notable distinctions. First, we extend the method to accommodate multiple pseudo observations with fractional weights, combining the approaches of weighted empirical likelihood (Glenn & Zhao, 2007) and an entropy balancing scheme (Hainmueller, 2012). Second, we pass to the limit, ensuring that the convex hull constraint is satisfied for all parameter values simultaneously. This resulting formulation naturally induces a form of regularization that removes the constraint. Our method’s main contributions encompass: (i) addressing the convex hull constraint for ETEL while retaining desirable asymptotic properties; (ii) enhancing stability and robustness of small-sample performance compared to existing methods; (iii) providing flexibility in Bayesian modelling and allowing one to incorporate a novel form of prior information.
This paper is organized as follows. In Section 2, we introduce the notation used in the paper and provide a brief overview of ETEL. Then, we propose a weighted version of ETEL that incorporates fractional pseudo-data with the maximum entropy reweighting scheme. In Section 3, we propose inducing regularization on the formulation of ETEL, exploring two equivalent approaches: (i) a limiting procedure with fractional pseudo-data and (ii) direct incorporation of a continuous exponential family distribution in the minimization of the KL divergence. We derive asymptotic properties of the proposed method. In Section 4, we evaluate the performance of the methods through simulation studies. In Section 5, we present an application to the estimation of median income for four-person families. Finally, we conclude with a discussion of directions for future research in Section 6. The proofs of the theoretical results are provided in the supplementary materials. The proposed method is implementation in the R package retel (Kim, 2024), available from the Comprehensive R Archive Network (CRAN) at https://cran.r-project.org/package=retel.
@@ 4.1 Posterior Coverage
To investigate the validity of RETEL for Bayesian inference, we begin by simulating a value of ${\theta}$ from a logistic distribution denoted as ${\textnormal{Logistic}(l,s)}$, where ${l}$ is the location parameter and ${s}$ is the scale parameter. Next, we generate ${n}$ observations from ${N(\theta,1)}$ and compute ${H}$ for the two versions of RETEL. Throughout the analysis, we employ ${p_{n}(\theta,\lambda)}$ in Equation 4 with ${\mu_{n,\theta}}={\overline{X}-\theta}$ and ${\Sigma_{n,\theta}}={1}$ for the univariate mean parameter ${\theta}$. For comparison purposes, we also compute ${H}$ using ETEL and AETEL. Keeping ${l}$ fixed at 00{0}, we repeat this procedure $10000$ times for each combination of ${n}\in{\{5,20,50,100\}}$, ${s}\in{\{1,5\}}$, and ${\tau_{n}}\in{\{1,\log n\}}$. We approximate the posterior distributions on a grid of ${\theta}$ values. Using the computed ${H}$ values, we conduct the Kolmogorov–Smirnov test to evaluate the uniformity of the distributions.
The resulting ${p}$-values are reported in Table 1, and Figure 3 displays the quantile-quantile plots for the distribution of ${H}$ versus ${U(0,1)}$ when ${n}={5}$, ${s}={5}$, and ${\tau_{n}}={1}$.
@@
The plots highlight the differences in the tails of the distributions that are not apparent from the ${p}$-values alone. With a smaller sample size of ${n}={5}$, RETEL tends to show a closer conformity to ${U(0,1)}$ compared to ETEL and AETEL. The impact of a larger prior variance (${s}={5}$) and a larger ${p_{c}(\theta)}$ (${\tau_{n}}={\log n}$) becomes more apparent when ${n}={50}$. As the sample size increases, the differences between the posterior distributions of the methods become negligible. All of the methods provide an excellent approximation to the null distribution when ${n}$ is 5050{50}50 or more. We emphasize that the Kolmogorov–Smirnov tests are based on a sample of $10000$ replicates and so are able to pick up quite small departures from the uniform distribution. Additional plots for the full results are provided in Section 8 in the supplementary materials.
Figure 3: Quantile-quantile plots for the distribution of ${H}$ versus ${U(0,1)}$ when ${n}={5}$, ${s}={5}$, and ${\tau_{n}}={1}$. The light-tailed distribution from ETEL is due to the convex hull constraint. AETEL produces a heavier-tailed distribution than the others.
@@
The results for ${l}={0}$ are presented in Table 2, where the prior mean matches the true parameter value. It can be seen that for all methods, as the sample size increases, the intervals become shorter and the coverage rates approach the target of ${95\%}$. As ${s}$ decreases, indicating stronger prior information on ${\theta}$ at 00{0}, higher coverage rates and shorter intervals are obtained. The differences between the methods are most pronounced when ${n}={5}$. The intervals obtained from ETEL exhibit significantly lower coverage rates compared to the other methods. AETEL produces the widest intervals with coverage rates higher than the nominal level. The wider intervals and departure from the nominal coverage rate are related to the boundedness problem of AEL, which arises due to the addition of one pseudo-observation (Emerson & Owen, 2009). On the other hand, RETEL yields coverage rates closer to the nominal level but features much shorter intervals compared to AETEL. Within RETEL, ${\textnormal{RETEL}_{f}}$ produces wider intervals with higher coverage rates than ${\textnormal{RETEL}_{r}}$, consistent with the findings from the plots in Figure 3.
@@ 4.2 Expected Kullback–Leibler Divergence
    Repeat Steps 1–3 $1000$ times and take the average of the estimates from Step 3.
Step 2 is implemented with two chains of length $5000$, ensuring that the potential scale reduction factor (Gelman & Rubin, 1992) remains below 1.11.1{1.1}1.1 on average for each method. For the regularized methods, ${\tau_{n}}={1}$ is used when ${n=2}$, and ${\tau_{n}}={\log n}$ is used otherwise. We implement EL using the R package melt (Kim et al., 2024).